# Survey of Natural Language Processing for Education: Taxonomy, Systematic Review, and Future Trends

### v3

@@ 3.4.2 Code Error Correction
Code Error Correction (CEC) fixes a buggy code snippet to make it error-free and more coherent. It is a challenging task as it requires the model to have a sufficient understanding of the code in long-distance dependency and make proper corrections to the erroneous code spans.
@@
Datasets. Defects4J [187], ManyBugs [188] and IntroClass [188] are some of the earlier datasets proposed for CEC tasks, focusing on Java and C languages respectively, but with small amounts of data. Lin et al. [189] proposed the first multi-lingual parallel corpus of CEC benchmarks, namely QuixBugs[189], which contains $4040$ buggy code snippets translated into both Java and Python. Xia et al. [190] combined the aforementioned Defects4J, ManyBugs, and QuixBugs datasets into a single CEC evaluation dataset to test the effect of PLM on the CEC task. Nowadays, the widespread use of collaborative software development platforms such as Github [^9] https://github.com/ makes it possible to collect and analyze code review data from open source projects. GitHub’s Forge provides a large amount of history-changing and bug-fixing commits from software projects, which can be the resource of CEC datasets in nature. Bugs2Fix [191] is one such dataset mined from GitHub bug-fixing commits, containing 2.3M bug-fix pairs (BFPS). CodeReview [192] is also a large-scale dataset of code changes and code reviews collected from open source projects, involving nine of the most popular programming languages. Because CodeReview has some low-quality code review data, and the data distribution in its test set is very similar to the data distribution in the pre-training and fine-tuning datasets, there is a potential risk when training with a model such as ChatGPT, which can affect the evaluation of the model’s generalization ability. Therefore, Guo et al. [193] create a new code review dataset CodeReview-New [193] based on it. By designing strict filtering rules to screen out low-quality data in CodeReview, they ended up retaining data from 467 repositories out of an initial 2,029.
Methods. The CEC and GEC tasks share similar principles and we can formulate both of them as sequence modeling, neural machine translation, and language modeling [102, 103]. Nevertheless, there are still some fundamental differences in their technical implementations. GEC tasks are typically solved by representing text as word embeddings or other forms of semantic vectors. While CEC tasks often require models to have a better understanding of long contexts, existing techniques capture the structural and semantic information of code by representing it as a sequence of symbols or an Abstract Syntax Tree (AST) [194]. Therefore, several pre-trained language models are trained using a large corpus of programming languages. CuBERT [167] is a pre-trained model for CEC tasks, which models only the Python language and is mainly applied to discriminative tasks of code understanding. CodeBERT [195] proposed in the same period is also modified on the basis of the BERT model, which is a multilingual programming language model that can handle dual-modal data (i.e. programming language (PL) and natural language (NL)). It supports modeling in 6 programming languages (i.e. Python, Java, JavaScript, PHP, Ruby, and Go) and combines pre-training tasks for replacement tag detection with hybrid objective functions, enabling the model to handle training using NL-PL data pairs. Both CuBERT and CodeBERT ignore the structural information of code. In response to this issue, Guo et al. [104] proposed the GraphCodeBert model to represent source code information based on AST data flow, so as to reflect deeper code information. As AST has a tree structure and cannot be directly used as the input of Transformer models, GraphCodeBERT allows the model to learn the AST data flow at the cost of losing a considerable part of information. In order to make more effective and comprehensive use of the code structure information provided by AST, UniXcoder [105] improves on the basis of GraphCodeBERT. They construct a one-to-one mapping function to convert an AST into a sequence structure, and then encode it together with the code text. In addition, some models use other ways rather than ASTs to better understand and adapt to the structural characteristics of a code snippet, and perform well. PLBART [106] adds additional denoising strategies to the BART model to correlate PL and NL text. It generates grammatically and logically valid code via large-scale sequence-to-sequence pre-training that incorporates the denoising strategies to better understand program grammars and logical flows. CodeT5 [107] uses the T5 model architecture and makes use of the identifier tagging information and comment information to better model the structural characteristics of code. CodePAD [108] designs an approach based on pushdown automaton (PDA), which takes into account the grammar constraints of programming languages. It ensures the availability of generated code and the correctness of grammar, and avoids the time-consuming and labor-intensive disadvantages of AST-based approaches.
@@ 4 Demonstration Systems
To facilitate applications on education scenarios, in this section, we present some off-the-shelf demo cases with brief illustrations of their usages for each task mentioned in Section 3. Information on these demos will be presented in Table V.
@@
For QA, OpenVINO [^10] https://docs.openvino.ai/2023.3/notebooks_section_2_model_demos.html is an open-source toolkit designed for diverse automatic QA, including table, visual, interactive, and LLMs-based QA. PrimeQA [196] is an open-source public repository designed to democratize QA research and simplify the replication of state-of-the-art methods, supporting the training of multilingual QA models with core functionalities like information retrieval, reading comprehension, and question generation.
@@
For MWP, MWPToolkit [197] offers a holistic and extensible framework that provides deep learning-based solvers, popular datasets, and a modular architecture that supports quick replication and innovation of MWP methods. IntMath [^11] https://www.intmath.com/ is an online math platform that combines a mathematical computation engine with artificial intelligence to parse and generate step-by-step natural language answers.
For QC, AutoQG [198] is an online service that provides multilingual question and answer generation, featuring a comprehensive toolkit for model fine-tuning, generation, and evaluation. Besides, Questimator [199] is a system that offers MCQs generation for educational assessment with a certain topic.
@@
For AES and ACS, LinggleWrite [200] is a writing coach tool that offers writing suggestions, proficiency assessments, grammatical error detection and corrective feedback. IntelliMetric [^12] https://www.intellimetric.com/direct/ is an online multilingual AES platform that supports writing evaluation by delivering instant scores with the accuracy and consistency of human experts. Qualified [^13] https://www.qualified.io/ and CoderPad [^14] https://coderpad.io/ are online platforms designed for coding assessments, offering interactive environments to evaluate coding skills.
For GEC and CEC, ALLECS [201] and Grammarly [^15] https://www.grammarly.com/ are web applications that provide online GEC service by offering lightweight GEC systems optimized for users with slow internet connections and comprehensive writing assistance, respectively. CodeGeeX [202] is an open-source and multilingual pre-trained model that enables syntax and function-correct code generation and can be integrated as IDE extensions.
Demo Task Form URL
OpenVINO QA toolkit https://docs.openvino.ai/2023.3/notebooks_section_2_model_demos.html
PrimeQA QA toolkit https://github.com/primeqa
MWPToolkit MWP toolkit https://github.com/LYH-YF/MWPToolkit
IntMath MWP platform https://www.intmath.com/
AutoQG QC application https://autoqg.net/
Questimator QC application https://github.com/methodical-software/questimator
LinggleWrite AES application https://f.linggle.com/
IntelliMetric AES platform https://www.intellimetric.com/direct/
Qualified ACS platform https://www.qualified.io/
CoderPad ACS platform https://coderpad.io/
ALLECS GEC application https://sterling8.d2.comp.nus.edu.sg/allecs/
Grammarly GEC application https://www.grammarly.com/
CodeGeeX CEC extension https://codegeex.cn/
TABLE V: Information on demos that can be applied in education domain.
5 Future Trends
According to the taxonomy and advanced techniques of NLP in the education domain, we discuss several promising future directions for different tasks.
More datasets in education domain. We have summarized the commonly used datasets for educational NLP, which range from question answering to error correction, we also observe there is still a demand for datasets fitting in more diverse education scenarios. Regarding QG tasks, instead of simply considering the single-turn question generation based on the context and answer, we may assume that multi-turn question generation will frequently happen in a tutoring scenario. Specifically, a leading question around a certain knowledge topic is important in guiding students to learn an ultimate subject goal. The distinction between such leading QG and traditional QG challenges lies in the connection between the leading generated questions and the final goal [11]. However, we find there is no existing challenge that has been defined in this way. Regarding AES and GEC tasks, event though we have collected a set of AES and GEC datasets for supporting the researches of natural language learning, there is a lack of AES and GEC datasets in non-English languages, which is fundamental in developing systems for low-resource languages. Furthermore, more interpretable datasets should be proposed to explicit more on the instructional feedback. For example, diverse dimensional explanation could be given in AES datasets and richer error types aligned to the specific grammatical knowledge schema could be provided in GEC datasets. Regarding ACS tasks, public ACS datasets are in high demand as the testbed for evaluating the ACS systems.
The controllable usage of LLMs. Due to the advancement of LLMs, recent studies have started to incorporate LLMs with various educational NLP tasks. Researchers have explored to prompt LLMs for solving zero/few-shot QA, generating questions over knowledge bases [203], and correcting grammatical errors [204] but there remain some drawbacks to be resolved. For example, in QA tasks, even though an LLM is able to act as a “decision center” to understand the questions as well as context and conduct parsing and reasoning, it may perform poorly in a specific knowledge domain. In this case, techniques like Retrieval-Augmented-Generation (RAG) and Multi-Agents are leveraged to enhance the performance of the LLM [33, 48]. But here comes some new questions. For example, the presence of noise or contradictory information retrieving from textbooks can detrimentally affect the quality of LLMs’ outputs. In QG tasks, strongly aligning the outputs of LLMs to the given textbook contexts and generating more controllable questions is an open question. Similarly, in GEC tasks, multiple studies [184] point out that applying LLMs on GEC tasks has a high risk of over-correction, that is failing to preserve the semantics of the original sentences. Therefore, we expect more studies could be proposed to investigate the controllable leverage of LLMs on GEC tasks.
Intervention of difficulty-level control. Most existing NLP techniques for education applications neglect difficulty-level control, which is an important factor in the procedure of learning, comprehension, and assessment at different levels. For example, in QG tasks, generating questions that could be aligned to the level of students is significant in teaching. Even though there are multiple studies [57, 58, 59] trying to model the difficulty level in the QG task, they simply define with the semantic or reasoning complexity of the questions and fail to align the level to the difficulty level of syllabus. For generating distractors, we rarely observe existing studies considering the difficulty level. But this is critical to identify a false candidate answer that is not only relevant to the question but also in accord with the difficulty level of students. In EC tasks, researchers focus more on the accuracy of the correction but neglect the level of the corrected sentences, including the level of substitute words, the syntactic and semantic level of the corrected sentences. Further researches should fuse the difficulty level with various tasks to make it serve educational NLP applications better.
Interpretable educational NLP. Existing studies are able to obtain good results on various tasks like question answering, question construction, automated assessment, and error correction. However, there are limited techniques developed with the perspective of interpretability. As shown in the technical review of QA tasks, recent studies try to explicit the process of thinking in solving Textbook QA and MWPs tasks [24, 127]. In other tasks, interpretability should also be taken into account. It would be promising to show a process of thinking path to make the construction of the questions more rational. For AS and EC tasks, by showing the fine-grained explanation to the scoring and correction, it would not only improve tutoring experience but also benefit the methods as well as systems to diagnose their functions. Even though, we notice a few studies start to develop interpretable GEC systems [174], interpretablity should be further studied and explored with more attention.
Methods with adaptive learning. Even though NLP applications can be applied to education domain, there are a few studies involving adaptive learning to different tasks. For QG task, we notice some researchers [62, 63] attempt to model the learning trajectory of students by incorporating their historical answers to predict their performance on the new questions, which are deemed as a significant feature to determine the difficulty level of the next question. This benefits to build a more customized education system. Such adaptive learning is in high demand to collaborate with other tasks. For example, in EC tasks, when we correct the sentences and codes, we can consider the level of the students based on their input to decide the way to correct. If it is a simple sentence with a mis-used word, we should not substitute it with an obscure words. Similarly, for a piece of erroneous code, we should not correct its errors using the knowledge beyond the knowledge of the students. In this case, developing methods with adaptive learning for various NLP tasks is important.
@@ 5 Future Trends
Integrated systems for education. Even though an individual task can be well-formulated and solved via specific techniques, there is a demand of integrated systems for education. Baladn et al. [205] tuned open-source LLMs for generating teacher responses in BEA 2023 Shared Task. EduChat [206] is a more general LLM in education domain, which is able to automatically assess essays, provide emotional support, and conduct Socratic teaching with the forms of chatting and question answering. In the future, more capabilities like answering textbook questions, automated quiz making, should be integrated to empower the system, making it perform as an efficient teaching assistant. Meanwhile, an integrated could ensemble the tutoring abilities in the dimension of different subjects such as math, computer science, linguistics and so on. In addition, exploring more interactive ways between the users and machines of an integrated system is worth studying. Generating responses from the system based on the question from the users may be not practical. In real-scenario, automatic answer or instruction prompting instead of waiting for the questions from students would help provide teaching assistance without raising up a question in purpose.
@@ 6 Conclusions
This survey attempted to provide a comprehensive overview of NLP in the education domain, as depicted in Figure 2. We highlight that NLP techniques can be applied to various procedure of education. After process and analysis the raw data from diverse data source, NLP applications can be directly serve to tutoring and learning. Our survey highlights four representative tasks (i.e., question answering, question construction, automated assessment, and error correction) with their fine-grained sub-tasks as the focus of this survey due to their significant impact in teaching and learning. According to the taxonomy, we go deep into these tasks and illustrate the evolution of their techniques. For each task, we describe its task definition, unique challenges in education domain, and its corresponding solutions. Early studies developed based on rule-based, deep learning-based methods to solve these tasks. Thank to the advancement of LLMs, a number of methods start collaborating LLMs with traditional methods. This innovation has led to a easier paradigm that improve the performance and decrease the demand of the annotation data. However, LLMs also encounter some new issues like domain knowledge incorporation and over-correction when we apply them to specific tasks. Therefore, future studies need to fuse LLMs to the diverse tasks with the consideration of solving the their drawbacks. Then, to help readers apply applications on education scenarios, we present the off-the-shelf toolkits and demonstrations. Despite the strides made in NLP applications on education domain, there still remains improvement room for diverse tasks. In the last section, we suggest some future directions for NLP in the education domain.
We would like to note that besides the highlighted tasks in this survey, there are some other tasks that can contribute to the educational scenarios. For example, knowledge graph construction [12, 207, 208] aims to organize and extract the knowledge points from the unstructured data sources to form a knowledge graph, which is able to demonstrate the connection between the schema for various subjects and denote to the downstream applications. Moreover, another NLP application controllable text generation [209, 210, 211] aims to generate text with given style strength, length, or components, which can help teachers prepare sentences teaching materials for the linguistics courses. In summary, the demand for educational applications is increasing and educational NLP will thrive to be a promising research area. We hope this survey will give a comprehensive picture of educational NLP and we encourage more contribution in this field.

### v4

@@ 3.4.2 Code Error Correction
Code Error Correction (CEC) fixes a buggy code snippet to make it error-free and more coherent. It is a challenging task as it requires the model to have a sufficient understanding of the code in long-distance dependency and make proper corrections to the erroneous code spans. It shares similar principles as GEC tasks and we can formulate both of them as neural machine translation [254, 255]:
$L_{\text{CEC}}=-\mathbb{E}_{(X,Y)\in\mathcal{D}}\log P_{\theta}(Y\mid X)$
Here, $X$ and $Y$ are the input and output code sequences.
@@
Datasets. Defects4J [256], ManyBugs [257] and IntroClass [257] are some of the earlier datasets proposed for CEC tasks, focusing on Java and C languages respectively, but with small amounts of data. Lin et al. [258] proposed the first multi-lingual parallel corpus of CEC benchmarks, namely QuixBugs[258], which contains $40$ buggy code snippets translated into both Java and Python. Xia et al. [259] combined the aforementioned Defects4J, ManyBugs, and QuixBugs datasets into a single CEC evaluation dataset to test the effect of PLM on the CEC task. Bugs2Fix [260] is a dataset mined from GitHub bug-fixing commits, containing 2.3M bug-fix pairs (BFPS). CodeReview [229] and CodeReview-New [230] are also large-scale datasets of code reviews collected from open-source projects, involving nine of the most popular programming languages.
Methods. Existing techniques capture the structural and semantic information of code by representing it as a sequence of symbols or an Abstract Syntax Tree (AST) [261], which involves some unique challenges for CEC tasks.
  • The structural information of code has been ignored. Both CuBERT [70] and CodeBERT [262] ignore the structural information of code. In response to this issue, Guo et al. [75] proposed the GraphCodeBert model to represent source code information based on AST data flow, so as to reflect deeper code information. GraphCodeBERT adapts Transformer and allows the model to learn the AST data flow. The follow-up UniXcoder [76] improves GraphCodeBERT by converting an AST into a sequence structure. PLBART [77] incorporates the denoising strategies to better understand program grammar and logical flows. CodeT5 [227] uses the T5 model architecture and makes use of the identifier tagging and comment information to better model the structural characteristics of code.
  • CEC systems have resource overhead issue. CEC systems’ high demands for memory and computing power may disrupt the user’s workflow and decrease productivity [79]. CodePAD [78] designs an approach based on pushdown automaton (PDA), which takes into account the grammar constraints of programming languages and reduces the unnecessary computational cost.
  • CEC systems encounter overfitting issue Overfitting refers to the issue where generated patches only pass the developer-written test suite but fail to generalize to other potential test cases [80]. Nilizadeh et al. [81] introduced a method that identifies correct patches based on the similarity of test case execution between the buggy and patched programs, which reduces overfitting by evaluating patches across a broader range of execution scenarios. Ye et al. [82] proposed an overfitting detection system, which leverages static code features, such as variable types, code structure, and contextual syntax information, to classify patches as overfitting or correct.
The performance and memory load of CEC systems could be improved more with the consideration of the informal and rigorous teaching environment for coding teaching.
@@ 4 Demonstration Systems
Task Demos
QA OpenVINO $\dagger$ PrimeQA $\dagger$ TableQAKit $\dagger$ PiggyBack $\ast$ ASCENT $\diamond$ UKP-SQUARE $\diamond$ LOCALRQA $\ast$
MWP MWPToolkit $\dagger$ IntMath $\diamond$ MWPRanker $\ast$
QC AutoQG $\ast$ SAQUET $\dagger$ AnswerQuest $\diamond$ Leaf $\ast$
AA LinggleWrite $\ast$ IntelliMetric $\diamond$ EXPATS $\dagger$ Qualified $\diamond$ CoderPad $\diamond$
GEC ALLECS $\diamond$ Grammarly $\diamond$ ERRANT $\dagger$ WAMP $\ast$ GECko+ $\ast$ Effidit $\diamond$
[…]
@@
To facilitate applications on education scenarios, in this section, we present some off-the-shelf demo cases with brief illustrations of their usage for each task mentioned in Section 3. Information on these demos will be presented in Table V. We annotate demonstrations with “platform”, “application”, “toolkit” and “extension”. Teachers and students could utilize platforms and extensions featured with interfaces and user instructions as AI helpers. Researchers could utilize toolkits and applications featured with development frameworks and flexible modules to implement new methods quickly.
@@
For QA, OpenVINO [^9] https://docs.openvino.ai/2023.3/notebooks_section_2_model_demos.html is an open-source toolkit designed for diverse automatic QA, including table, visual, interactive, and LLMs-based QA. PrimeQA [263] is an open-source public repository designed to democratize QA research and simplify the replication of state-of-the-art methods, supporting the training of multilingual QA models with core functionalities like information retrieval, reading comprehension, and question generation. TableQAKit [264] is an open-source toolkit designed for table-based QA, offering a unified platform with various datasets, methods, and LLM integrations. PiggyBack [265] is a user-friendly VQA platform that simplifies the use of visual-language pre-trained models with a browser-based interface and comprehensive task support. Besides, other QA demos [266, 267, 268] are also available for both educators and researchers.
@@
For MWP, MWPToolkit [269] offers a holistic and extensible framework that provides deep learning-based solvers, popular datasets, and a modular architecture that supports quick replication and innovation of MWP methods. IntMath [^10] https://www.intmath.com/ is an online math platform that combines a mathematical computation engine with artificial intelligence to parse and generates step-by-step natural language answers. MWPRanker [270] is a tool to retrieve similar math word problems based on expression tree similarity.
For QC, AutoQG [271] is a service that provides multilingual question-and-answer generation, featuring a comprehensive toolkit for model fine-tuning, generation, and evaluation. SAQUET [272] proposes a toolkit to assess the structural and pedagogical quality of MCQs. AnswerQuest [273] is a system that integrates QA and QG to create Q&A items for enhancing reading comprehension of multi-paragraph documents. Besides, other QC demos [274] are also available for both educators and researchers.
@@
For AES and ACS, LinggleWrite [275] is a writing coach tool that offers writing suggestions, proficiency assessments, grammatical error detection, and corrective feedback. IntelliMetric [^11] https://www.intellimetric.com/direct/ is an online multilingual AES platform that supports writing evaluation by delivering instant scores with the accuracy and consistency of human experts. EXPATS [276] is an open-source framework for AES, enabling rapid experimentation with various models, while integrating visualization and interpretability tools. Qualified [^12] https://www.qualified.io/ and CoderPad [^13] https://coderpad.io/ are online platforms designed for coding assessments, offering interactive environments to evaluate coding skills.
For GEC and CEC, ALLECS [277] and Grammarly [^14] https://www.grammarly.com/ are web applications that provide online GEC service by offering lightweight GEC systems optimized for users with slow internet connections and comprehensive writing assistance, respectively. Errant [278] is a toolkit that extracts and classifies edits from parallel sentences, enabling error type evaluation and dataset standardization. WAMP [279] is a web-based annotation tool designed to efficiently create annotated corpora for GEC, with features like customizable error tags and file export for system evaluation. Other demos [280, 281] can also be applied for both researchers and educators. CodeGeeX [282] is an open-source and multilingual pre-trained model that enables syntax and function-correct code generation and can be integrated as IDE extensions. CodeHelp [283] is an LLM-powered assistant that provides real-time programming guidance and explanations, helping students solve coding issues independently without giving direct solutions. MarsCode [284] is an LLM-based multi-agent framework for automated bug fixing.
5 Challenges and Future Trends
In this section, we will discuss the challenges and future trends when applying NLP technology in real educational scenarios.
Generalization over subjects and languages. Even though there is a range of datasets from question answering to error correction, we also observe there is still a demand for datasets and techniques fitting in more subjects and languages. For example, for AES and GEC tasks, there is a lack of AES and GEC datasets in non-English languages, which is vital for language education in non-English countries. For CEC tasks, we found limited datasets for Python language, which is a widely used programming language today. Similarly, applications on different subjects are vital in extending the NLP techniques to education. Even though there is a series of datasets covering the subjects of grades 6-8 or K-12, for higher education, which includes some advanced subjects like philosophy, and history, we still face a shortage of data. In the future, a comprehensive dataset is desired to cover subjects of different learning stages and different tasks for both teaching and research.
@@ 5 Challenges and Future Trends
Deployed LLM-based systems for education. Even though an individual task can be well-formulated and solved via specific techniques, there is a demand for integrated systems for education. Baladn et al. [285] tuned open-source LLMs for generating teacher responses in BEA 2023 Shared Task. EduChat [286] is a more general LLM in the education domain, which is able to automatically assess essays, provide emotional support, and conduct Socratic teaching with the forms of chatting and question answering. In the future, an integrated LLM-based tutoring system should be featured with stronger capabilities like answering textbook questions, and automated quiz-making in the dimension of different subjects such as math, computer science, linguistics, and so on. Regarding system deployment, although open-source NLP technology demos, including toolkits, online platforms, and applications, have been widely developed (see Section 4). The limited hardware and resources make it challenging to apply advanced NLP in education, prompting studies focused on improving feasibility in this area. For example, studies [287, 288] of MWP solving have shown that applying knowledge distillation can significantly reduce model size and improve inference speed while maintaining high solution accuracy. Further optimization is needed considering the edge computing and efficiency.
Adaptive learning for teaching and learning. Even though NLP applications can be applied to the education domain, there are a few studies involving adaptive learning. In the QG task, we notice some researchers [43, 44] attempt to model the learning trajectories of students by incorporating their historical answers to predict their performance on the new questions, which are deemed as a significant feature to determine the difficulty level of the next question. This helps to build a more customized education system. Such adaptive learning is in high demand to collaborate with other tasks. For example, in EC tasks, when we correct the sentences and codes, we can consider the learning trajectory of the students to decide the way to correct them. If it is a simple sentence with a misused word, we should not substitute it with an obscure word. Besides modeling the learning trajectories, the intervention of difficulty-level control mechanisms should be integrated into the education system. Even though there are multiple studies [34, 35, 36] trying to model the difficulty level in the QG task, they fail to align the level to the difficulty level of the syllabus.
Interpretability for education. Existing studies are able to obtain good results on various tasks like question answering, question construction, automated assessment, and error correction. However, there are limited techniques developed with the perspective of interpretability. As shown in the technical review of QA tasks, recent studies try to explicate the process of thinking in solving TQA and MWPs tasks [289, 27]. In other tasks, interpretability should also be taken into account. It would be promising to show a thinking path and make the construction of the questions more rational. For AS and EC tasks, showing a fine-grained explanation of the scoring and correction would not only improve the tutoring experience but also benefit the methods as well as systems to diagnose their functions. Even though we notice a few studies have started to develop interpretable GEC systems [204], interpretability should be further studied and explored with more attention.
Ethical considerations of NLP techniques. Applying NLP techniques for education raises various ethical issues, such as fairness, safety, privacy, and dependency. Li et al. [290] systematically introduce the fairness of LMs. As noted by Chinta et al. [291], racial bias in AI-driven essay scoring systems has resulted in darker-skinned students receiving lower scores than other students for essays of comparable quality. Regarding privacy concerns, when students use third-party generative AI tools for assignments or reflections, they may unintentionally expose sensitive academic data, risking privacy law violations. These tools lack transparency around how data is stored, processed, or reused [292]. To address this issue, researchers [293, 294] have shown that federated learning protects student privacy by training models locally on sensitive data and sharing only model updates, enabling collaborative model development without exposing raw data. For dependency, Krupp et al. [295] indicate an overreliance on ChatGPT among prospective physics teachers and students, affecting their use of real-world contextualization and leading to poorer results in complex tasks. Multiple education-related studies also attempt to reduce educators’ or students’ overreliance on current techniques [283, 296].
@@ 6 Conclusions
This survey attempted to provide a comprehensive overview of NLP in the education domain. We highlight that NLP techniques can be applied to various procedures of education. Our survey highlights four representative tasks (i.e., question answering, question construction, automated assessment, and error correction) with their fine-grained sub-tasks as the focus of this survey due to their significant impact on teaching and learning. According to the taxonomy, we go deep into these tasks and illustrate the evolution of their techniques. Despite the strides made in NLP applications in the education domain, there still remains room for the improvement of diverse tasks. In the last section, we suggest some future directions for NLP in the education domain.
We would like to note that besides the highlighted tasks in this survey, there are other studies that can contribute to the educational scenarios. For example, knowledge graph construction [84] aims to organize and extract the knowledge points from the unstructured data sources to form a knowledge graph, which is able to demonstrate the connection between the schema for various subjects and support downstream applications. In summary, the demand for educational applications is increasing and educational NLP will thrive to be a promising research area. We hope this survey will give a comprehensive picture of educational NLP and we encourage more contribution in this field.